OpenStudy (ajprincess):

plzzz help Show that if A and B are two non-empty sets, having n elements in common then AXB anb BXA have n^2 elements in common.

OpenStudy (anonymous):

IF N^2 elements be in common so should be A=B

OpenStudy (anonymous):

cause of (x,y)=(y,x) then x=y

myininaya (myininaya):

One sec still working on it

OpenStudy (ajprincess):

ha k thanx

myininaya (myininaya):

brb getting tea

myininaya (myininaya):

@zarkon I need more brains

myininaya (myininaya):

\[\text{ If } |A \cap B|=1, \text{ then there is element } x \text{ such that } x \in A \cap B \] \[(A \times B ) \cap( B \times A)=\{(x,x)\} => |A \cap B|=1=1^2\]

myininaya (myininaya):

\[\text{ If } | A \cap B| =k , \text{ there there are elements} x_1,x_2,x_3,...x_k \in A \cap B \]

OpenStudy (anonymous):

well i can't put my opinion here cause i can't understand the question or i didn't study that Branch yet So lead the Ship @myininaya & i'm following u :)

myininaya (myininaya):

\[(A \times B ) \cap (B \times A)= \{ (x_1,x_1), (x_1,x_2),(x_1,x_3),...(x_1,x_k),(x_2,x_1),...\] \[(x_2,x_k),...(x_k,x_1),...(x_k,x_k) \}\]

myininaya (myininaya):

Do you see that is k^2 elements like we have all the k elements with (x_1,blah) then we have k elements with (x_2,blah) then k elements with (x_3,blah) then .... k elements with (x_k, blah)

myininaya (myininaya):

you have k, k times

myininaya (myininaya):